Find the frequency of the sound wave in the tube of the length L = 1m, if it’s both ends are (a) open; (b) closed, (c) one end is opened and another end is closed. The speed of sound v =340 m/s.


1
Expert's answer
2021-03-24T16:08:59-0400

Length of tube L=1m


(a) When Both ends are open-

"f=\dfrac{v}{2l}=\dfrac{340}{2(1)}=170 Hz"


(b) when Both ends are closed-

"f=\dfrac{v}{4l}=\dfrac{340}{4(1)}=85Hz"


(c) When one end is open and one is closed-

"f=\dfrac{3v}{4l}=\dfrac{3\times 340}{4(1)}=255 Hz"